# ”Living Within Four Walls”: Exploring Emotional and Social Dynamics in Mobile Usage During Home Confinement

### v3

@@ Abstract.
Home confinement, a situation experienced by individuals for reasons ranging from medical quarantines, rehabilitation needs, disability accommodations, and remote working, is a common yet impactful aspect of modern life. While essential in various scenarios, confinement within the home environment can profoundly influence mental well-being and digital device usage. Using the COVID-19 lockdown as a case study, this research explores the emotional and social effects of prolonged home confinement on mobile device usage. We conducted an in-situ study with 32 participants, analyzing three weeks of mobile usage data to assess emotional well-being and social dynamics in restricted environments. Our findings reveal that app usage patterns serve as strong indicators of emotional states, offering insights into how digital interactions can reflect and influence well-being during isolation. This study highlights the potential for developing targeted interventions and support systems for individuals in long-term home confinement, including those with chronic illness, recovery needs, or permanent remote work situations.
Emotion, Home Confinement, Mobile Sensing, COVID-19 Lockdown, Social Roles, Well-Being
copyright: acmcopyrightjournalyear: 2024doi: 10.1145/1122445.XXXXccs: Human-centered computing Ubiquitous and mobile computingccs: Applied computing
1. Introduction
@@
Home confinement, defined as the restriction of movements within the home (52) due to illness, recovery, disability, or other personal circumstances, is a common phenomenon that everyone may experience during their lives (38; 11). Research has shown that such confinement can adversely affect mental well-being, leading to stress, anxiety, and depression (6; 66). For instance, Bartoszek et al. (6) found that quarantined individuals experienced heightened levels feelings of depression, insomnia, loneliness, and fatigue. Understanding these impacts is essential for developing effective targeted interventions to support individuals during such periods.
@@
Traditionally, mental well-being during home confinement has been assessed using survey-based instruments (66; 22), which can be time-consuming, labour-intensive, and prone to biases such as social desirability and reflection bias (23; 59). Recent advances in mobile sensing, however, offer a more objective, unobtrusive and real-time alternative (61; 15). While methods relying on physical movement data (e.g., GPS and accelerometers) have been widely used for mental wellbeing modelling such as social anxiety, affect, depression and social isolation (61; 15), they are less effective in the context of indoor confinement, where movement patterns (e.g., step counts, visited locations) are limited (61). This study shifts focus to more granular data, such as app usage patterns, to provide deeper insights into emotional well-being during confinement.
@@
The diversity of individual circumstances complicates the study of home confinement. Factors such as the reason for confinement, personal health conditions, and social environment can vary greatly, complicating the creation of a standardized model. However, the COVID-19 lockdown provided a unique, global and natural context in which the majority of individuals experienced similar restrictions, with a marked increased reliance on digital technologies (57; 70). This shared experience allows for a more uniform analysis of how mobile device usage reflects emotional states, social roles, and work-life dynamics during such periods. Espeically, while this study uses the lockdown as a descriptive case study (87), the insights gained may also be applicable to other forms of home confinement, such as those experienced by individuals with chronic illnesses, in post-recovery rehabilitation, or those adapting to permanent remote work arrangements.
@@
In this study, we explore the relationship between mobile device interactions, emotional well-being, and social roles, during COVID-19 lockdown, as a case study for home confinement. We focus on two key aspects of well-being: emotions and social roles, as both are critical to mental health (17; 54). By utilizing the Valence and Arousal dimensions from the Circumplex Model of Affect (67) and Social Role from the Role Theory (8), we aim to provide a comprehensive understanding of how mobile device usage during lockdown reflects emotional and social dynamics. Our research seeks to address the following questions: RQ1. How are emotions and social roles during COVID-19 home confinement perceived through self-reports and inferred from app usage patterns? RQ2. Can mobile usage data be used to model emotional well-being during home confinement, and which features are most predictive?
@@
To answer these questions, we conducted a three-week data collection during the COVID-19 lockdown in Melbourne, Australia, involving 32 participants. We analyzed perceived emotional states, social roles, and work-life balance through Experience Sampling Method (ESM) (7) and end-of-day surveys. We also proposed a novel method, App-based Information Gain Temporal Segmentation (AIGTS), to extract user activities from app usage records by identifying segments and transition times between apps during the day. The resulting insights offer a deeper understanding of how app usage during confinement reflects emotional well-being and social dynamics, with implications for developing digital interventions to support individuals in long-term home confinement. All scripts for this study are open source. [^1] Link will be shared after the double-blind review. In summary, our contributions are as follows:
@@
  • We collected a rich dataset from 32 participants during a 3-week COVID-19 home confinement period in Melbourne, Australia. This dataset includes 502,485 records of users smartphone and desktop usage data, 1,749 ESM responses, and 265 end-of-day surveys. It provides a unique, real-world snapshot of emotional well-being and social dynamics during lockdown.
@@
  • We introduced the App-based Information Gain Temporal Segmentation (AIGTS) method, an unsupervised technique for extracting user activities (e.g., video-watching, socializing, deep work) from app usage data. This method autonomously identified 16 distinct user activities, providing granular insights into how digital behaviours relate to emotional well-being.
@@
2. Related work
2.1. Impact of the Home Confinement on Mental Well-Being
@@
Home confinement, whether due to edical conditions, disability, or societal events like pandemics, can significantly affect mental well-being. Prolonged isolation, particularly when coupled with limited physical activity and disrupted social interactions, is linked to increased risks of anxiety, depression, stress, and loneliness (76; 14; 10). The psychological impact is often compounded by the loss of routine and control, leading to emotional disturbances such as irritability, confusion, and anger (64; 39). These effects may persist even after restrictions are lifted (31).
@@
The COVID-19 pandemic underscored these effects on a global scale. The World Health Organization (84) reported that the psychological toll of lockdowns and social distancing measures often surpassed the physical health risks. Studies during the pandemic have documented a significant increases in mental health issues, with depression and anxiety becoming widespread (85; 16; 37). For example, Khubchandani et al. (36) found depression (39%) and anxiety (42%) surged in the U.S. post-lockdown. Similarly, Fountoulakis et al. (20) reported increased anxiety (45%) and suicidal thoughts (10.4%) in Greece. Loneliness was also a major issue, especially among those in long-term isolation (37).
@@
Despite extensive research on the mental well-being impacts of various forms of home confinement, much of it is geographically limited or reflects different phases of isolation. This variability makes it challenging to fully understand the broader, long-term psychological effects of home confinement in diverse contexts. To address this gap, RQ1 examines how emotions and social roles are experienced during COVID-19 home confinement, using self-reports and app usage patterns to infer these dynamics.
2.2. Digital Device Usage During Home Confinement
@@
During periods of home confinement, digital devices play a crucial role in maintaining communication, accessing information, and fulfilling work and social roles. With restricted physical mobility, devices such as smartphones, tablets, and computers become essential for staying connected to social networks and managing day-to-day activities (32). For instance, Ratan et al. (62) found that, during home confinement, smartphones became a vital tool for work-related communication and personal connections, often replacing face-to-face interactions.
@@
While many studies have explored digital device usage in general contexts, the COVID-19 lockdown provided a unique opportunity to study digital usage behaviour under prolonged confinement. With workplaces and schools shifting to remote environments, the demand for digital communication tools surged (79; 19; 5). Video conferencing platforms, social media, and messaging apps saw massive increases in usage (70; 57), and overall digital device use skyrocketed. A global survey found 70% of internet users reported increased smartphone usage during the lockdown (70), emphasizing how integral smartphones became in managing both work and personal life. Similarly, Majumdar et al. (45) observed that office workers spent significantly more time on digital devices, including smartphones, desktops, and televisions, during the lockdown. Jonnatan et al. (33) also found that mobile device usage rose during the pandemic, with urban participants reporting higher use and greater concerns about overuse, addiction, and negative health impacts.
@@
In the post-pandemic era, many individuals continue to rely heavily on digital devices due to the persistence of remote work and online communication (43; 29). Studies suggest that behaviours such as prolonged screen time and device overuse, which emerged during the pandemic, have continued into the post-lockdown phase (46). This highlights the importance of understanding digital device usage patterns in today’s technology-reliant environments, particularly in contexts of home confinement.
@@
2.3. Inferring Mental Well-Being Using Mobile Sensing Technologies
@@
Advances in mobile sensing technologies have opened new avenues for understanding human mental well-being, including mood (49; 88; 48), engagement (24; 28; 2), and mental health issues like depression and anxiety (15; 61; 81). Many studies have integrated data from multiple sensors to improve prediction accuracy. For example, Gao et al. (24) used wearable and environmental sensors to predict student engagement, while Wang et al. (81) tracked depression using passive mobile sensing and wearables. Jaques et al. (30) modelled student mood using physiological signals and mobile behaviour.
@@
While multi-source sensing enhances accuracy, there is increasing interest in using mobile sensing data alone, given its convenience and ability to capture a range of behaviours like physical activity, communication patterns, and location (25). Studies have shown that mobile usage data can predict psychological states such as anxiety (61; 77), mood (88; 48), and emotion (75). For instance, Zhang et al. (88) developed an app to detect compound emotions from smartphone usage, and Tlachac et al. (77) used call and text logs to predict anxiety and depression during the pandemic.
@@
Research on app sequences has also provided insights into daily behaviours, such as social roles (3) and interruptibility (55; 56). Anderson et al. (3) demonstrated that app sequences correlate with individuals’ work and private roles, while Okoshi et al. (55; 56) used app sequences to predict activity breakpoints for intelligent notification delivery. However, most studies have focused on identifying breakpoints rather than modeling user activity or predicting contextual variables. Additionally, earlier studies often relied on basic aggregation methods, which fail to capture transitions between activities. Thus, RQ1 also explores emotions and social roles by analyzing app usage patterns.
@@
While mobile sensing technologies have shown potential in predicting well-being, gaps remain in understanding how these tools can be applied specifically during home confinement. Home confinement introduces unique behaviours—such as increased reliance on digital devices and altered perceptions of time—that differ from daily life. Although some studies (77; 53; 74) have examined psychological states during the pandemic, they focus on broader environments rather than the specific context of home confinement. Understanding this unique setting could provide valuable insights for other confinement scenarios, such as post-operative recovery, parental leave, or aging. Therefore, RQ2 investigates how mobile sensing can model emotions during home confinement and identify key features for prediction.
3. Study Overview
@@
We conducted a three-week field study in Melbourne, Australia, from 24 August 2020 to 13 September 2020, during the strictest Stage IV COVID-19 restrictions. This study use Melbourne’s experience as a lens to explore broader home confinement scenarios, such as those due to medical conditions, recovery needs, or remote work. The following section details the background, participants, and data collected.
3.1. A Case Study of COVID-19 Lockdown in Melbourne, Australia
Figure 1. The periods of six lockdowns and our data collection
@@ 3.1. A Case Study of COVID-19 Lockdown in Melbourne, Australia
To prevent the community spread of COVID-19, Melbourne has been under six lockdowns since March 2020, with a total of 263 days of lockdowns, making it the city with the longest COVID-19 lockdown in the world (47; 73). Our data collection was conducted during the second lockdown period with Stage IV restrictions (the strictest measures), which lasted from 9 July 2020 to 27 October 2020, with our study conducted from 24 August 2020 to 13 September 2020 (see Figure 1). Among the six lockdown periods, this was one of the longest and strictest lockdown periods, during which a state of disaster was declared on 2 August 2020, followed by Stage IV restrictions in metropolitan Melbourne (51).
@@
Under the Stage IV restrictions in force at the time of our data collection, a daily curfew was implemented from 8 PM to 5 AM, forbidding anyone from leaving their homes except for special reasons. Residents were only allowed to shop and exercise within 5 km of their homes, and all students returned to home-based learning. Other restrictions that were previously applied only to specific districts were extended to the entire Melbourne metropolitan area. On 13 September 2020, several restrictions were eased in Melbourne, including reduced curfews and loosened rules on outdoor exercise and social interactions. During home confinement due to the lockdown, people showed a strong reliance on technology, particularly for school and communication (62; 70).
3.2. Procedures

### v4

# Exploring Emotional and Social Dynamics in Mobile Usage During Home Confinement

@@ Abstract
Home confinement, a situation experienced by individuals for reasons ranging from medical quarantines, rehabilitation needs, disability accommodations, and remote working, is a common yet impactful aspect of modern life. While essential in various scenarios, it can profoundly influence mental well-being and digital device usage. Using the COVID-19 lockdown as a case study, this research explores the emotional and social effects of home confinement on mobile device usage. We conducted an in-situ study with 32 participants, analyzing three weeks of mobile usage data to assess emotional well-being and social dynamics in restricted environments. We found that app usage patterns serve as strong indicators of emotional states, offering insights into how digital interactions can reflect and influence well-being during isolation. This study highlights the potential for developing interventions and support systems for individuals in long-term home confinement, including those with chronic illness, recovery needs, or permanent remote work situations.
keywords:
Human-centred Computing , Mobile Sensing , Mental wellbeing , Affective computing , Home confinement , Emotion
journal: Computers in Human Behavior
[…]
[aff3]organization=RMIT University, city=Melbourne, postcode=3000, country=Australia
[…]
[aff_g]organization=University of Kassel, city=Kassel, postcode=34131, country=Germany
1 Introduction
@@
Home confinement, defined as the restriction of movements within the home [1] due to illness, recovery, disability, or other personal circumstances, is a common phenomenon that everyone may experience during their lives [2, 3]. Research has shown that such confinement can adversely affect mental well-being, leading to stress, anxiety, and depression [4, 5]. For instance, Bartoszek et al. [4] found that quarantined individuals experienced heightened levels feelings of depression, insomnia, loneliness, and fatigue. Understanding these impacts is essential for developing effective targeted interventions to support individuals during such periods.
@@
Traditionally, mental well-being during home confinement has been assessed using survey-based instruments [5, 6], which can be time-consuming, labour-intensive, and prone to biases such as social desirability and reflection bias [7, 8]. Recent advances in mobile sensing, however, offer a more objective, unobtrusive and real-time alternative [9, 10]. While methods relying on physical movement data (e.g., GPS and accelerometers) have been widely used for mental wellbeing modelling such as social anxiety, affect, depression and social isolation [9, 10], they are less effective in the context of indoor confinement, where movement patterns (e.g., step counts, visited locations) are limited [9]. This study shifts focus to more granular data, such as app usage patterns, to provide deeper insights into emotional well-being during confinement.
@@
The diversity of individual circumstances complicates the study of home confinement. Factors such as the reason for confinement, personal health conditions, and social environment can vary greatly, complicating the creation of a standardized model. However, the COVID-19 lockdown provided a unique, global and natural context in which the majority of individuals experienced similar restrictions, with a marked increased reliance on digital technologies [11, 12]. This shared experience allows for a more uniform analysis of how mobile device usage reflects emotional states, social roles, and work-life dynamics during such periods. Especially, while this study uses the lockdown as a descriptive case study [13], the insights gained may also be applicable to other forms of home confinement, such as those experienced by individuals with chronic illnesses, in post-recovery rehabilitation, or those adapting to permanent remote work arrangements.
@@
In this study, we explore the relationship between mobile device interactions, emotional well-being, and social roles during home confinement, using data collected during the COVID-19 lockdown as a case study. We focus on two key aspects of well-being: emotions and social roles, as both are critical to mental health [14, 15]. By utilizing the Valence and Arousal dimensions from the Circumplex Model of Affect [16] and Social Role from the Role Theory [17], we aim to provide a comprehensive understanding of how mobile device usage during lockdown reflects emotional and social dynamics. Our research seeks to address the following questions: RQ1. How are emotions and social roles during home confinement perceived through self-reports and inferred from app usage patterns? RQ2. Can mobile usage data be used to model emotional well-being during home confinement, and which features are most predictive?
@@
To answer these questions, we conducted a three-week data collection during the COVID-19 lockdown in Melbourne, Australia, involving 32 participants. We analyzed perceived emotional states, social roles, and work-life balance through Experience Sampling Method (ESM) [18] and end-of-day surveys. We also proposed a novel method, App-based Information Gain Temporal Segmentation (AIGTS), to extract user activities from app usage records by identifying segments and transition times between apps during the day. The resulting insights offer a deeper understanding of how app usage during confinement reflects emotional well-being and social dynamics, with implications for developing digital interventions to support individuals in long-term home confinement.
All scripts for this study are open source. [^4] Link will be shared after the double-blind review. In summary, our contributions are as follows:
  • 1.
@@
    We collected a rich dataset from 32 participants during a 3-week COVID-19 home confinement period in Melbourne, Australia. This dataset includes 502,485 records of users’ smartphone and desktop usage data, 1,749 ESM responses, and 265 end-of-day surveys. It provides a unique, real-world snapshot of emotional well-being and social dynamics during lockdown.
  • 2.
@@
  • 3.
    We introduced the App-based Information Gain Temporal Segmentation (AIGTS) method, an unsupervised technique for extracting user activities (e.g., video-watching, socializing, deep work) from app usage data. This method autonomously identified 16 distinct user activities, providing granular insights into how digital behaviors relate to emotional well-being.
  • 4.
@@
2 Related work
2.1 Impact of the Home Confinement on Mental Well-Being
@@
Home confinement, whether due to medical conditions, disability, or societal events like pandemics, can significantly affect mental well-being. Prolonged isolation, particularly when coupled with limited physical activity and disrupted social interactions, is linked to increased risks of anxiety, depression, stress, and loneliness [19, 20, 21]. The psychological impact is often compounded by the loss of routine and control, leading to emotional disturbances such as irritability, confusion, and anger [22, 23]. These effects may persist even after restrictions are lifted [24].
@@
The COVID-19 pandemic underscored these effects on a global scale. The World Health Organization [25] reported that the psychological toll of lockdowns and social distancing measures often surpassed the physical health risks. Studies during the pandemic have documented a significant increases in mental health issues, with depression and anxiety becoming widespread [26, 27, 28]. For example, Khubchandani et al. [29] found depression (39%) and anxiety (42%) surged in the U.S. post-lockdown. Similarly, Fountoulakis et al. [30] reported increased anxiety (45%) and suicidal thoughts (10.4%) in Greece. Loneliness was also a major issue, especially among those in long-term isolation [28].
@@
Despite extensive research on the mental well-being impacts of various forms of home confinement, much of it is geographically limited or reflects different phases of isolation. This variability makes it challenging to fully understand the broader, long-term psychological effects of home confinement in diverse contexts. To address this gap, RQ1 examines how emotions and social roles are experienced during home confinement, using self-reports and app usage patterns to infer these dynamics.
2.2 Digital Device Usage During Home Confinement
@@
During periods of home confinement, digital devices play a crucial role in maintaining communication, accessing information, and fulfilling work and social roles. With restricted physical mobility, devices such as smartphones, tablets, and computers become essential for staying connected to social networks and managing day-to-day activities [31]. For instance, Ratan et al. [32] found that, during home confinement, smartphones became a vital tool for work-related communication and personal connections, often replacing face-to-face interactions.
@@
While many studies have explored digital device usage in general contexts, the COVID-19 lockdown provided a unique opportunity to study digital usage behavior under prolonged confinement. With workplaces and schools shifting to remote environments, the demand for digital communication tools surged [33, 34, 35]. Video conferencing platforms, social media, and messaging apps saw massive increases in usage [12, 11], and overall digital device use skyrocketed. A global survey found 70% of internet users reported increased smartphone usage during the lockdown [12], emphasizing how integral smartphones became in managing both work and personal life. Similarly, Majumdar et al. [36] observed that office workers spent significantly more time on digital devices, including smartphones, desktops, and televisions, during the lockdown. Jonnatan et al. [37] also found that mobile device usage rose during the pandemic, with urban participants reporting higher use and greater concerns about overuse, addiction, and negative health impacts.
@@
In the post-pandemic era, many individuals continue to rely heavily on digital devices due to the persistence of remote work and online communication [38, 39]. Studies suggest that behaviors such as prolonged screen time and device overuse, which emerged during the pandemic, have continued into the post-lockdown phase [40]. This highlights the importance of understanding digital device usage patterns in today’s technology-reliant environments, particularly in contexts of home confinement.
@@
2.3 Inferring Mental Well-Being Using Mobile Sensing Technologies
@@
Advances in mobile sensing technologies have opened new avenues for understanding human mental well-being, including mood [41, 42, 43], engagement [44, 45, 46], and mental health issues like depression and anxiety [10, 9, 47]. Many studies have integrated data from multiple sensors to improve prediction accuracy. For example, Gao et al. [44] used wearable and environmental sensors to predict student engagement, while Wang et al. [47] tracked depression using passive mobile sensing and wearables. Jaques et al. [48] modeled student mood using physiological signals and mobile behavior.
@@
While multi-source sensing enhances accuracy, there is increasing interest in using mobile sensing data alone, given its convenience and ability to capture a range of behaviors like physical activity, communication patterns, and location [49]. Studies have shown that mobile usage data can predict psychological states such as anxiety [9, 50], mood [42, 43], and emotion [51]. For instance, Zhang et al. [42] developed an app to detect compound emotions from smartphone usage, and Tlachac et al. [50] used call and text logs to predict anxiety and depression during the pandemic.
@@
Research on app sequences has also provided insights into daily behaviors, such as social roles [52] and interruptibility [53, 54]. Anderson et al. [52] demonstrated that app sequences correlate with individuals’ work and private roles, while Okoshi et al. [53, 54] used app sequences to predict activity breakpoints for intelligent notification delivery. However, most studies have focused on identifying breakpoints rather than modeling user activity or predicting contextual variables. Additionally, earlier studies often relied on basic aggregation methods, which fail to capture transitions between activities. Thus, RQ1 also explores emotions and social roles by analyzing app usage patterns.
@@
While mobile sensing technologies have shown potential in predicting well-being, gaps remain in understanding how these tools can be applied specifically during home confinement. Home confinement introduces unique behaviors, such as increased reliance on digital devices and altered perceptions of time, that differ from daily life. Although some studies [50, 55, 56] have examined psychological states during the pandemic, they focus on broader environments rather than the specific context of home confinement. Understanding this unique setting could provide valuable insights for other confinement scenarios, such as post-operative recovery, parental leave, or aging. Therefore, RQ2 investigates how mobile sensing can model emotions during home confinement and identify key features for prediction.
3 Study Overview
@@
We conducted a three-week field study in Melbourne, Australia, from 24 August 2020 to 13 September 2020, during the strictest Stage IV COVID-19 restrictions. This study uses Melbourne’s experience as a lens to explore broader home confinement scenarios, such as those due to medical conditions, recovery needs, or remote work. The following section details the background, participants, and data collected.
3.1 A Case Study of COVID-19 Lockdown in Melbourne, Australia
Figure 1: The periods of six lockdowns and our data collection
@@ 3.1 A Case Study of COVID-19 Lockdown in Melbourne, Australia
To prevent the community spread of COVID-19, Melbourne has been under six lockdowns since March 2020, with a total of 263 days of lockdowns, making it the city with the longest COVID-19 lockdown in the world [57, 58]. Our data collection was conducted during the second lockdown period with Stage IV restrictions (the strictest measures), which lasted from 9 July 2020 to 27 October 2020, with our study conducted from 24 August 2020 to 13 September 2020 (see Figure 1). Among the six lockdown periods, this was one of the longest and strictest lockdown periods, during which a state of disaster was declared on 2 August 2020, followed by Stage IV restrictions in metropolitan Melbourne [59].
@@
Under the Stage IV restrictions in force at the time of our data collection, a daily curfew was implemented from 8 PM to 5 AM, forbidding anyone from leaving their homes except for special reasons. Residents were only allowed to shop and exercise within 5 km of their homes, and all students returned to home-based learning. Other restrictions that were previously applied only to specific districts were extended to the entire Melbourne metropolitan area. On 13 September 2020, several restrictions were eased in Melbourne, including reduced curfews and loosened rules on outdoor exercise and social interactions. During home confinement due to the lockdown, people showed a strong reliance on technology, particularly for school and communication [32, 12].
3.2 Procedures